# Classical and quantum thermodynamics in a non–equilibrium regime: Application to Stirling engine

### v1

@@ III Numerical Demonstration
Although the results presented in Tables 1 and  2 hold for any non–equilibrium system consisting of subsystem and bath, the extensive variables and waste heat that appear in these relationships are system-specific, and there is no general theory on the basis of which non–equilibrium thermodynamic potentials can be obtained. However, it is possible to evaluate them as functions of the intensive and extensive variables using an optimization algorithm. Here, as a demonstration, we evaluate non–equilibrium thermodynamic potentials numerically using the thermodynamic SB model.Koyanagi and Tanimura (2024a)
III.1 Thermodynamic system–bath model
We employ the Ullersma–Caldeira–Leggett (or Brownian) model,Ullersma (1966a, b); Caldeira and Leggett (1983); Grabert, Schramm, and Ingold (1988); Weiss (2012); Tanimura and Wolynes (1991, 1992) which is ideal for thermodynamic simulations because the subsystem and bath are well defined, and rigorous numerical solutions can be obtained in both classical and quantum cases under any time-dependent external perturbation. Many of the favorable features for thermodynamic investigations arise from the presence of a counter term, which allows us to include the contribution of the SB interaction in the bath.Tanimura (2015, 2006, 2020); Tanimura and Wolynes (1991, 1992); Kato and Tanimura (2013) By introducing multiple heat baths at different temperatures controlled by time-dependent SB coupling functions, we can investigate isothermal, isentropic, thermostatic, and entropic processes. The total Hamiltonian for isothermal and thermostatic processes is written asKoyanagi and Tanimura (2024a)
@@ III.1 Thermodynamic system–bath model
is the unperturbed Hamiltonian of the subsystem with mass $m$ and potential $U(\hat{q})$ described by momentum $\hat{p}$ and position $\hat{q}$. The internal energy is then evaluated as $U_{\rm A}^{\rm neq}(t)=\mathrm{tr}\{\hat{H}_{\rm A}^{0}\hat{\rho}_{\rm A}(t)\}$, where $\hat{\rho}_{\rm A}(t)$ is the reduced density operator of the subsystem. The external perturbation is expressed as $\hat{H}_{\rm A}^{\prime}(t)\equiv-x(t)\hat{X}_{\rm A}$, where $\hat{X}_{\rm A}$ is an operator of the subsystem coordinate [i.e., $\hat{X}_{\rm A}(\hat{q})$] and $x(t)$ is regarded as the thermodynamic intensive variable. The extensive variable is evaluated as $X_{\rm A}^{\rm neq}(t)=\mathrm{tr}\{\hat{X}_{\rm A}\hat{\rho}_{\rm A}(t)\}$.
Although the conventional SB model has been limited to the investigation of isothermal processes at constant temperature, we can extend it to describe thermostatic processes in which temperature varies with time by introducing $N$ independent heat baths, each in the thermal equilibrium state at the inverse temperature $\beta_{k}\equiv 1/k_{\rm B}T_{k}$ connected to or disconnected from subsystem A using the window function $\xi_{k}(t)$.Koyanagi and Tanimura (2024a) The $k$th bath Hamiltonian is expressed as an ensemble of harmonic oscillators and is given by
@@
where the momentum, position, mass, and frequency of the $j$th bath oscillator are given by $\hat{p}_{j}^{k}$, $\hat{x}_{j}^{k}$, $m_{j}^{k}$, and $\omega_{j}^{k}$, respectively. Here, we consider the situation where $N$ independent heat baths, each in the thermal equilibrium state $\exp(-\beta_{k}\hat{H}_{\rm I+B}^{k})$ at the inverse temperature $\beta_{k}\equiv 1/k_{\rm B}T_{k}$, are connected or disconnected to subsystem A according to a control function $\xi_{k}(t)$. The bath temperature is effectively expressed as
@@
or the inverse temperature as $\beta(t)=[k_{\rm B}T(t)]^{-1}$. As a spectral distribution function for the $k$th bath $J^{k}(\omega)\equiv\sum_{j}{\hbar(c_{j}^{k}A_{k})^{2}}/({2m_{j}^{k})\omega_{j}^{k}}\delta(\omega-\omega_{j}^{k})$, we consider the Ohmic case described by
@@
and assume that the time scale of the quantum thermal fluctuations $\beta(t)\hbar/2\pi$ is shorter than the time scale of the external perturbations $\beta(t)$ and $x(t)$. This allows us to extend the low-temperature quantum Fokker–Planck equations (LT-QFPE)Ikeda and Tanimura (2019) for the set of Wigner distribution functions $W_{\vec{n}}(p,q;t)$, where ${\vec{n}}$ is the index of hierarchy member, and the Kramers equation for phase space distribution function $W(p,q;t)$ to apply the thermostatic case by introducing time-dependent Matsubara frequencies $\nu(t)=1/\hbar\beta(t)$. Expressions for these equations are presented in Ref. Koyanagi and Tanimura, 2024a. The source code for thermodynamic HEOM in phase space representation and thermodynamic Kramers equation are provided as a separated paper.Koyanagi and Tanimura (2024b).
It should be noted that in the non-Ohmic case, the temperature may change even on the time scale on which the noise correlations are defined. This makes it difficult to apply the fluctuation–dissipation theorem to characterize the noise, and thus a simple replacement $\beta\rightarrow\beta(t)$ is not allowed.
@@ III.2 DL work and extensive variables
We express the solution for the reduced density elements under any $x(t)$ in the Wigner representation using the zeroth member of the hierarchical Wigner functions as $W(p,q,t)\equiv W_{\vec{0}}(p,q;t)$. In the classical limit $\hbar\rightarrow 0$, ${W}(q,p;t)$ corresponds to the classical distribution function. We use this to define the change in DL intensive work over time corresponding to power and heat flow as follows:
@@
From the definition (2) of extensive work and the TDL transformations (6) and (7), the extensive heat current is evaluated as
@@
III.3 Evaluation of the non–equilibrium Planck potential
Figure 1: Schematic of the evaluation of the non–equilibrium Planck potential $\Xi_{\rm A}^{\rm neq}(t)$ at $t_{k}$ and $t_{k+1}$. A non–equilibrium process driven by the intensive variables $\beta(t)$ and $\tilde{x}(t)$ is described as the non–equilibrium distribution $W(p,q;t)$ expressed by the black arrows. The equilibrium states for different sets of $\beta^{\rm qst}_{n_{k}}$ and $\tilde{x}_{n_{k}}^{\rm qst}$ are expressed by the red dashed line. For each state $k$ to the equilibrium state $n_{k}$, we minimize $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}+(\Xi_{\rm A}^{\rm qst})_{n_{k}}$ by choosing the trial functions $\beta_{k}^{\rm trial}(t^{\prime})$ and $\tilde{x}_{k}^{\rm trial}(t^{\prime})$, where $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}$ is the DL minimum intensive work for $k\rightarrow n_{k}$ and $(\Xi_{\rm A}^{\rm qst})_{n_{k}}$ is the quasi-static Planck potential at the state $n_{k}$. We can then evaluate the non–equilibrium Planck potential at the state $k$ as $(\Xi_{\rm A}^{\rm neq})_{k}=(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}+(\Xi_{\rm A}^{\rm neq})_{n_{k}}$.
From Eq. (51), we obtain the expression for the non–equilibrium Planck potential $(\Xi_{\rm A}^{\rm neq})_{k}=\Xi_{\rm A}^{\rm neq}[\beta(t_{k}),\tilde{x}(t_{k})]$ of the non–equilibrium-to-equilibrium process $k\rightarrow n_{k}$ as
@@ III.3 Evaluation of the non–equilibrium Planck potential
where $n_{k}$ represents an equilibrium state, $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}$ is the DL minimum intensive work for $k\rightarrow n_{k}$, and $(\Xi_{\rm A}^{\rm qst})_{n_{k}}$ is the quasi-static Planck potential at state $n_{k}$. As shown in Eq. (55), because the right-hand side of Eq. (40) is independent of the choice of $n_{k}$, to evaluate $(\Xi_{\rm A}^{\rm neq})_{k}$, we do not specify $n_{k}$ to minimize $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}+(\Xi_{\rm A}^{\rm qst})_{n_{k}}$.
To perform numerical calculations, we express Eq. (40) with Eq. (15) in terms of trial functions $\beta^{\rm trial}(t_{k},t^{\prime})$ and $\tilde{x}^{\rm trial}(t_{k},t^{\prime})$ as
$Xi_{\rm A}^{\rm target}[\beta^{\rm trial}(t),x^{\rm trial}(t)]=\Xi^{\rm qst}_{\rm A}(\beta^{\rm qst}_{n_{k}},x^{\rm qst}_{n_{k}})$
@@
where $\Xi_{\rm A}^{\rm target}[\beta^{\rm trial}(t),\tilde{x}^{\rm trial}(t)]$ is the target function to be minimized and $\Xi_{\rm A}^{\rm qst}(\beta_{n_{k}}^{\rm qst},\tilde{x}_{n_{k}}^{\rm qst})$ is the quasi-static Planck potential at $n_{k}$, which is also evaluated from the LT-QFPE and Kramers equation. The DL trail function $\tilde{x}_{k}^{\rm trial}(t)$ in Eq.(III.3) is evaluated from $x_{k}^{\rm trial}(t)$ as $\tilde{x}_{k}^{\rm trial}(t)=\beta_{k}^{\rm trial}(t)x_{k}^{\rm trial}(t)$. To reduce the computational cost, we assume that the optimized trial functions become constant after the characteristic time for the equilibration $\Delta t$, as $\beta^{\rm qst}_{n_{k}}=\beta_{k}^{\rm trial}(t_{k}+\Delta t)$ and $x_{n_{k}}^{\rm qst}=x_{k}^{\rm trial}(t_{k}+\Delta t)$.
As the trial functions for time $t^{\prime}>t_{k}$, we chose the ${N_{\beta}}$th and ${N_{x}}$th Taylor expansion forms expressed as
@@
where $\beta_{k}^{(n)}$ and $x_{k}^{(n)}$ are the $n$th-order Taylor coefficients. Thus, the functional minimization of $\Xi_{\rm A}^{\rm target}[\beta^{\rm trial}(t),\tilde{x}^{\rm trial}(t)]$ becomes a multivariable function minimization for $\beta_{k}^{(n)}$ and $x_{k}^{(n)}$.
Then, $\Xi_{\rm A}^{\rm neq}(t)$ is evaluated as follows (see Fig. 1). First, we perform the non–equilibrium simulations for given $\beta(t)$ and $\tilde{x}(t)$ to obtain $W_{\vec{n}}(p,q;t)$ for each $t_{k}$. The minimum work from the non–equilibrium state $k$ to the equilibrium state $n_{k}$ expressed as $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}$ is then evaluated using an optimization algorithm for $\beta_{k}^{\rm trial}(t^{\prime})$ and $\tilde{x}_{k}^{\rm trial}(t^{\prime})$. From Eq. (40), we set this value as $(\Xi_{\rm A}^{\rm neq})_{k}$. By repeating this procedure for different values of $k$, we obtain $\Xi_{\rm A}^{\rm neq}(t_{k})$ at each step. The other potentials can be evaluated from the TDL transformations.
III.4 Quantum and classical Stirling engine
The non–equilibrium thermodynamic potentials are state variables as functions of the non–equilibrium intensive and extensive variables, which are also state variables, while $Q_{\rm A}^{\rm wst}$ is not, and, in the quasi-static limit, they reduce to conventional thermodynamic potentials. These non–equilibrium potentials are useful because they can be used to analyze thermodynamic processes using work diagrams, as in the case of equilibrium thermodynamics. By finding a thermodynamic process to minimize waste heat, we can construct a heat machine with maximum efficiency under non–equilibrium conditions, although the efficiency is lower than the Carnot limit.Koyanagi and Tanimura (2022a, b)
Here, we demonstrate how to evaluate non–equilibrium potentials for a case in which $\beta(t)$ and $x(t)$ are specified. For this purpose, we consider the classical/quantum Stirling engine (CSE/QSE) consisting of four steps: (i) a hot isothermal process, (ii) a transition from a hot to cold thermostatic process, (iii) a cold isothermal process, and (iv) a transition from a cold to hot thermostatic process for an anharmonic potential system. Since the purpose of the calculation is to illustrate how the potentials are calculated, we fix the changes in $\beta(t)$ and $x(t)$ as shown in Fig. 2 with $\tau=20$ and do not optimize them to improve their efficiency.

### v2

@@ III Numerical Demonstration
Although the results presented in Tables 1 and  2 hold for any non-equilibrium system consisting of subsystem and bath, the extensive variables and entropy production that appear in these relationships are system-specific, and there is no general theory on the basis of which non-equilibrium thermodynamic potentials can be obtained. However, it is possible to evaluate them as functions of intensive and extensive variables using an optimization algorithm. Here, as a demonstration, we evaluate non-equilibrium thermodynamic potentials numerically using the thermodynamic SB model.Koyanagi and Tanimura (2024b, c)
III.1 Thermodynamic system-bath model
We employed the Ullersma–Caldeira–Leggett (or Brownian) model,Ullersma (1966a, b); Caldeira and Leggett (1983); Grabert, Schramm, and Ingold (1988); Weiss (2012); Tanimura and Wolynes (1991, 1992) which is ideal for thermodynamic simulations because the subsystem and bath are well defined, and rigorous numerical solutions can be obtained in both classical and quantum cases under any time-dependent external perturbation. Many of the favorable features for thermodynamic investigations arise from the presence of a counter term, which allows us to include the contribution of the SB interactions in the bath.Tanimura (2015, 2006, 2020); Tanimura and Wolynes (1991, 1992); Kato and Tanimura (2013)
By introducing multiple heat baths at different temperatures controlled by time-dependent SB coupling functions, we can investigate isothermal, isentropic, thermostatic, and entropic processes. The total Hamiltonian for isothermal and thermostatic processes is written asKoyanagi and Tanimura (2024b, c)
@@ III.1 Thermodynamic system-bath model
is the unperturbed Hamiltonian of a subsystem with mass $m$ and potential $U(\hat{q})$ described by momentum $\hat{p}$ and position $\hat{q}$. The internal energy is then evaluated as $U_{\rm A}^{\rm neq}(t)=\mathrm{tr}\{\hat{H}_{\rm A}^{0}\hat{\rho}_{\rm A}(t)\}$, where $\hat{\rho}_{\rm A}(t)$ is the reduced density operator of the subsystem. The external perturbation is expressed as $\hat{H}_{\rm A}^{\prime}(t)\equiv-x(t)\hat{X}_{\rm A}$, where $\hat{X}_{\rm A}$ is an operator of the subsystem coordinate [i.e., $\hat{X}_{\rm A}(\hat{q})$], and $x(t)$ is the thermodynamic intensive variable. The extensive variable is evaluated as $X_{\rm A}^{\rm neq}(t)=\mathrm{tr}\{\hat{X}_{\rm A}\hat{\rho}_{\rm A}(t)\}$.
Although the conventional SB model has been limited to the investigation of isothermal processes at constant temperature, we can extend it to describe thermostatic processes in which temperature varies with time by introducing $N$ independent heat baths, each in the thermal equilibrium state at the inverse temperature $\beta_{k}\equiv 1/k_{\rm B}T_{k}$ connected to or disconnected from subsystem A using the window function $\xi_{k}(t)$.Koyanagi and Tanimura (2024b) The $k$th bath Hamiltonian is expressed as an ensemble of harmonic oscillators and is given by
@@
where the momentum, position, mass, and frequency of the $j$th bath oscillator are given by $\hat{p}_{j}^{k}$, $\hat{x}_{j}^{k}$, $m_{j}^{k}$, and $\omega_{j}^{k}$, respectively. Here, we consider the situation, in which $N$ independent heat baths, each in the thermal equilibrium state $\exp(-\beta_{k}\hat{H}_{\rm I+B}^{k})$ at the inverse temperature $\beta_{k}\equiv 1/k_{\rm B}T_{k}$, are connected or disconnected to subsystem A according to a control function $\xi_{k}(t)$. The bath temperature can be effectively expressed as
@@
or the inverse temperature as $\beta(t)=[k_{\rm B}T(t)]^{-1}$. As a spectral distribution function of the $k$th bath $J^{k}(\omega)\equiv\sum_{j}{\hbar(c_{j}^{k}A_{k})^{2}}/({2m_{j}^{k})\omega_{j}^{k}}\delta(\omega-\omega_{j}^{k})$, we consider the Ohmic case described by
@@
and assume that the time scale of quantum thermal fluctuations $\beta(t)\hbar/2\pi$ is shorter than the time scale of the external perturbations $\beta(t)$ and $x(t)$. This allows us to extend the low-temperature quantum Fokker–Planck equations (LT-QFPE)Ikeda and Tanimura (2019) for the set of Wigner distribution functions $W_{\vec{n}}(p,q;t)$, where ${\vec{n}}$ is the index of hierarchy members, and the Kramers equation for the phase space distribution function $W(p,q;t)$ to the thermostatic case by introducing time-dependent Matsubara frequencies $\nu(t)=1/\hbar\beta(t)$. The expressions for the thermodynamic quantum Fokker-Planck equations (T-QFPE) and thermodynamic Kramers equation (T-KE) are given in Refs. Koyanagi and Tanimura, 2024b, c. The source codes for them are also provided in the supplementary material.Koyanagi and Tanimura (2024c)
While the HEOM approach to thermodynamics treats the SB interaction as part of the main system in spin-boson systems,Kato and Tanimura (2015, 2016); Sakamoto and Tanimura (2021, 2020); Koyanagi and Tanimura (2022a, b, 2024a) this Brownian model treats it as part of a heat bath, including the counter term.Koyanagi and Tanimura (2024b, c) Assuming the Ohmic SDF, in the classical and high-temperature semi-classical cases, the dynamics described by this model exhibit a Markovian future, which can be treated in the framework of stochastic thermodynamics, while in the fully quantum case, it is not only non-Markovian, but also non-factorized owing to bathentanglement. In other words, the difference between the classical and quantum results represents a deviation from Markovian thermodynamics that arises when dealing with fully quantum processes.
It should be noted that for non-Markovian processes, the SDF-based description of a thermal bath breaks down when the correlation time of the noise is longer than the time scale of the bath temperature change.Koyanagi and Tanimura (2024c) In such cases, prescriptions that simply replace $\beta$ with $\beta(t)$ are not allowed, and the hierarchy members for each bath at different temperatures must be treated separately.Koyanagi and Tanimura (2024a) In this way, several baths can be operated simultaneously, although this is computationally expensive.Kato and Tanimura (2015, 2016)
@@ III.2 DL work and extensive variables
We express the solution for reduced density elements under any $x(t)$ in the Wigner representation using the zeroth member of the hierarchical Wigner functions as $W(p,q,t)\equiv W_{\vec{0}}(p,q;t)$. In the classical limit $\hbar\rightarrow 0$, ${W}(q,p;t)$ corresponds to the classical distribution function. We use this to define the change in DL intensive work over time, which corresponds to power and heat flow as follows:
@@
From the definition (2) of extensive work and the TDL transformations (6) and (7), the extensive heat current can be evaluated as
@@
III.3 Evaluation of the non-equilibrium Planck potential
Figure 1: Schematic of the evaluation of the non-equilibrium Planck potential $\Xi_{\rm A}^{\rm neq}(t)$ at $t_{k}$ and $t_{k+1}$. A non-equilibrium process driven by the intensive variables $\beta(t)$ and $\tilde{x}(t)$ is described as the non-equilibrium distribution $W(p,q;t)$ expressed by the black arrows. The equilibrium states for different sets of $\beta^{\rm qst}_{n_{k}}$ and $\tilde{x}_{n_{k}}^{\rm qst}$ are expressed by the red dashed line. For each state $k$ to the equilibrium state $n_{k}$, we minimize $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}+(\Xi_{\rm A}^{\rm qst})_{n_{k}}$ by choosing the trial functions $\beta_{k}^{\rm trial}(t^{\prime})$ and $\tilde{x}_{k}^{\rm trial}(t^{\prime})$, where $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}$ is the DL minimum intensive work for $k\rightarrow n_{k}$ and $(\Xi_{\rm A}^{\rm qst})_{n_{k}}$ is the quasi-static Planck potential at the state $n_{k}$. We can then evaluate the non-equilibrium Planck potential at state $k$ as $(\Xi_{\rm A}^{\rm neq})_{k}=(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}+(\Xi_{\rm A}^{\rm neq})_{n_{k}}$.
From Eq. (51), we obtain the following expression for the non-equilibrium Planck potential $(\Xi_{\rm A}^{\rm neq})_{k}=\Xi_{\rm A}^{\rm neq}[\beta(t_{k}),\tilde{x}(t_{k})]$ of the non-equilibrium-to-equilibrium process $k\rightarrow n_{k}$ as
@@ III.3 Evaluation of the non-equilibrium Planck potential
where $n_{k}$ represents the equilibrium state, $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}$ is the DL minimum intensive work for $k\rightarrow n_{k}$, and $(\Xi_{\rm A}^{\rm qst})_{n_{k}}$ is the quasi-static Planck potential at state $n_{k}$. As shown in Eq. (55), the right-hand side of Eq. (40) is independent of the choice of $n_{k}$; thus, to evaluate $(\Xi_{\rm A}^{\rm neq})_{k}$, we do not specify $n_{k}$ to minimize $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}+(\Xi_{\rm A}^{\rm qst})_{n_{k}}$.
To perform numerical calculations, we express Eq. (40) with Eq. (15) in terms of trial functionals $\beta^{\rm trial}(t_{k},t^{\prime})$ and $\tilde{x}^{\rm trial}(t_{k},t^{\prime})$ as
$X_{\rm A}^{\rm target}[\beta^{\rm trial}(t),x^{\rm trial}(t)]=\Xi^{\rm qst}_{\rm A}(\beta^{\rm qst}_{n_{k}},x^{\rm qst}_{n_{k}})$
@@
where $\Xi_{\rm A}^{\rm target}[\beta^{\rm trial}(t),\tilde{x}^{\rm trial}(t)]$ is the target function to be minimized and $\Xi_{\rm A}^{\rm qst}(\beta_{n_{k}}^{\rm qst},\tilde{x}_{n_{k}}^{\rm qst})$ is the quasi-static Planck potential at $n_{k}$, which is also evaluated from the T-QFPE and T-KE. The DL trial functional $\tilde{x}_{k}^{\rm trial}(t)$ in Eq. (III.3) is evaluated from $x_{k}^{\rm trial}(t)$ as $\tilde{x}_{k}^{\rm trial}(t)=\beta_{k}^{\rm trial}(t)x_{k}^{\rm trial}(t)$. To reduce computational cost, we assume that the optimized trial functionals are constant after the characteristic time for the equilibration $\Delta t$, as $\beta_{k}^{\rm trial}(t_{k}+\Delta t)=\beta^{\rm qst}_{n_{k}}$ and $x_{k}^{\rm trial}(t_{k}+\Delta t)=x_{n_{k}}^{\rm qst}$.
As trial functionals for time $t^{\prime}>t_{k}$, we chose the ${N_{\beta}}$th and ${N_{x}}$th Taylor expansion forms expressed as
@@
where $\beta_{k}^{(n)}$ and $x_{k}^{(n)}$ are the $n$th-order Taylor coefficients. Thus, the functional minimization of $\Xi_{\rm A}^{\rm target}[\beta^{\rm trial}(t),\tilde{x}^{\rm trial}(t)]$ becomes a multivariable functional minimization for $\beta_{k}^{(n)}$ and $x_{k}^{(n)}$.
Then, $\Xi_{\rm A}^{\rm neq}(t)$ is evaluated as follows (see Fig. 1). First, we perform the non-equilibrium simulations for given $\beta(t)$ and $\tilde{x}(t)$ to obtain $W_{\vec{n}}(p,q;t)$ for each $t_{k}$. The minimum work from the non-equilibrium state $k$ to the equilibrium state $n_{k}$ expressed as $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}$ is then evaluated by an optimization algorithm for $\beta_{k}^{\rm trial}(t^{\prime})$ and $\tilde{x}_{k}^{\rm trial}(t^{\prime})$. From Eq. (40), we set this value as $(\Xi_{\rm A}^{\rm neq})_{k}$. By repeating this procedure for different $k$ values, we obtain $\Xi_{\rm A}^{\rm neq}(t_{k})$ at each step. The other potentials can be evaluated by using the TDL transformations.
III.4 Thermostatic quantum and classical Stirling engines
The non-equilibrium thermodynamic potentials are state variables as functions of the non-equilibrium intensive and extensive variables, which are also state variables, whereas $Q_{\rm A}^{\rm wst}$ is not; in the quasi-static limit, they reduce to conventional thermodynamic potentials. These non-equilibrium potentials are useful because they can be used to analyze thermodynamic processes using work diagrams, as in the case of equilibrium thermodynamics. By identifyuing a thermodynamic process that minimizes entropy production, we can construct a heat machine with maximum efficiency under non-equilibrium conditions, although the efficiency is lower than the Carnot limit.Koyanagi and Tanimura (2022a, b)
Here, we demonstrate how to evaluate non-equilibrium potentials for the case in which $\beta(t)$ and $x(t)$ are specified. For this purpose, we consider a thermostatic classical/quantum Stirling engineKoyanagi and Tanimura (2024c) consisting of four steps: (i) a hot isothermal process, (ii) a transition from a hot-to-cold thermostatic process, (iii) a cold isothermal process, and (iv) a transition from a cold-to-hot thermostatic process for an anharmonic potential system. Since the purpose of the calculation is to illustrate how potentials are calculated, we fix the changes in $\beta(t)$ and $x(t)$ as shown in Fig. 2 with $\tau=20$ and do not optimize them to improve their efficiency.